OpenStudy (calculusxy):

A total of $20,000 is invested in two bonds that pay 8% and 9.5% in interest. The annual interest is $1675. How much is invested in each bond?

OpenStudy (calculusxy):

@Deeezzzz

OpenStudy (mertsj):

x = amount invested at 8% 20000-x= amount invested at 9.5%

OpenStudy (calculusxy):

How does 2000-x equal to 9.5%?

OpenStudy (mertsj):

it's 20000-x

OpenStudy (calculusxy):

Yeah sorry that.

OpenStudy (mertsj):

.08x = interest earned on 8% investment .095(20000-x)= interest earned on 9.5% investment

OpenStudy (mertsj):

Total interest earned is $1675 Write and solve the equation that says the sum of the two interest amounts is1675

OpenStudy (calculusxy):

.08x + .095(20000-x) = $1675

OpenStudy (mertsj):

Solve it.

OpenStudy (calculusxy):

0.08x + 0.095(20000-x) = 1675 0.08x + 1900 + 0.095x = 1675 0.175x + 1900 = 1675 0.175x = -375 2142?

OpenStudy (mertsj):

-.095x

OpenStudy (calculusxy):

For what? Why is it negative?

OpenStudy (mertsj):

\[.095(20000-x)=1900-.095x\]

OpenStudy (calculusxy):

oh yeah

OpenStudy (calculusxy):

0.705x = 1675 x = 2375

OpenStudy (mertsj):

\[.08x+.095(20000-x)=1675\] \[.08x+1900-.095x=1675\] \[-.015x+1900=1675\] \[-.015x=-225\] \[\frac{-.015x}{-.015}=\frac{-225}{-.015}\] \[x=15000\]

OpenStudy (calculusxy):

I am sorry. I am really tired so I can't focus as usual. I have to get it on track. :)

OpenStudy (mertsj):

So 15000 was invested at 8% 20000-x=5000 = amount invested at 9.5%

OpenStudy (calculusxy):

x= 15000

OpenStudy (calculusxy):

So those are the answers right?